OpenStudy (anonymous):

Bricklayers use the formula N = 7LH to estimate the number of bricks N needed to build a wall of height H and length L. What is the height of a wall that is 30 feet long and that requires 2,310 bricks to build?

OpenStudy (texaschic101):

N = 7LH N = 2310 L = 30 now sub 2310 = 7(30) * H 2310 = 210H 2310/210 = H 11 = H check.. N = 7LH 2310 = 7 * 30 * 11 2310 = 210 * 11 2310 = 2310 (correct) Height is 11 ft any questions ?